# arXiv:2601.10941v2[math.AP] 01 Mar 2026

### v2

@@ 3.2. General action estimates
We first note that for every $\lambda\in(\lambda_{n},\lambda_{n+1})$ and $u\in Q_{\lambda}$, we have
@@ 4. Further estimates in the case of a metric measure space
Next we observe that, since $u\in Q_{\lambda}$, the function $u$ is an eigenfunction of the operator $A-V_{u}$ corresponding to the eigenvalue $\lambda$. Here, we note that $A-V_{u}$ is a self adjoint operator defined as a form sum of the operators $A$ and $V_{u}$. Here we note that, by Hölder’s inequality, $V_{u}$ is a well-defined and continuous multiplication operator $L^{p}(\Omega)\to L^{p^{\prime}}(\Omega)$. Therefore, by (38) and duality, it also defines a compact operator $E\to E^{\prime}$. In other words, $V_{u}$ is a symmetric and relatively form compact perturbation of the operator $A$, which implies that $A-V_{u}$ is also bounded from below and has a compact resolvent, see [23, Chapter XIII.4 and Problem XIII.39]. Therefore $A-V_{u}$ admits a sequence of eigenvalues
@@ 5.2. Weyl’s law, large spectral gaps and large $L^{\infty}$ norms
$V_{u}\in L^{\infty}(\Omega;\mathbb{R}),\qquad V_{u}(x)=\left\{\begin{aligned} &\frac{f(u(x))}{u(x)},&&\qquad\text{$u(x)\neq 0$,}\\ &0,&&\qquad\text{$u(x)=0$}\end{aligned}\right$.
@@
(50) $\lambda_{n}(u)=\inf_{\stackrel{{\scriptstyle E^{\prime}\subset E}}{{\dim E^{\prime}\geq n}}}\sup_{v\in E^{\prime}\setminus\{0\}}\frac{1}{\|v\|_{H}^{2}}\Bigl(\langle A^{1/2}v,A^{1/2}v\rangle-\int_{\Omega}V_{u}v^{2}\,\mathrm{d}x\Bigr).$
@@
$\frac{1}{\|v\|_{H}^{2}}\Bigl|\int_{\Omega}V_{u}v^{2}\,\mathrm{d}x\Bigr|\leq\|V_{u}\|_{L^{\infty}(\mathcal{G})}\qquad\text{for every $u\in E$, $v\in E\setminus\{0\}$,}$
@@
$\lambda_{n+1}-\lambda\leq\lambda_{n+1}-\lambda_{n+1}(u)\leq\|V_{u}\|_{L^{\infty}(\mathcal{G})}\leq\tilde{g}(\|u\|_{L^{\infty}(\mathcal{G})}),$
which proves the claim. ∎
@@ 5.3. ODE arguments and the $L^{\infty}$ to $L^{2}$ estimate
for all $M\geq R$. Since $\frac{1}{\kappa_{0}}R\geq s_{1}$, Lemma 5.7 implies that $\tau_{\lambda,M}\leq\ell/2$ for all $M\geq R$. Hence, the solution $u_{\mid e}$ does at least one full period on $e$. Thus, there exists an interval $I\subseteq e$ with $|I|\geq|e|/2\geq\ell/2$ on which $e$ is periodic. Lemma 5.5 implies that
@@ 5.4. Sign of solutions in the presence of a constant-sign eigenfunction
Taking scalar products with $\xi$, we obtain
$(Au,\xi)_{L^{2}}=\lambda\int_{\mathcal{G}}u\xi\,\mathrm{d}x+\int_{\mathcal{G}}f(u)\xi\,\mathrm{d}x.$
Since $\xi$ is an eigenvalue of $A$, we have that
$(Au,\xi)_{L^{2}}=\lambda_{i}\int_{\mathcal{G}}u\xi\,\mathrm{d}x.$
@@
$(\lambda_{i}-\lambda)\int_{\mathcal{G}}u\xi\,\mathrm{d}x=\int_{\mathcal{G}}f(u)\xi\,\mathrm{d}x,$
@@
As already pointed out in the introduction, the first eigenfunction is positive in $\mathcal{G}$ in the case of continuity-Kirchhoff conditions. For a further study of conditions ensuring positivity of the first eigenfunction, one may refer to [18, Section 4.5].
@@ 6. The nonlinear biharmonic equation on the $2$-torus
denotes the sums of two squares function. A result from analytic number theory (see [15, Chapter 16.9]) implies that
@@
where $d(j)$ is the number of divisors of $j$. From this inequality, one can deduce (see [15, Theorem 315]) that for every $\varepsilon>0$ there exists $C_{\varepsilon}>0$ with
@@ Acknowledgements
D.G. is a Francqui Fellow of the Belgian American Educational Foundation at Brown University and was a Research Fellow of the F.R.S.-FNRS when this research was initiated. He would like to thank the Francqui Foundation, the Belgian American Educational Foundation, the University of Mons, the F.R.S.-FNRS and the Goethe-Universität Frankfurt for supporting this research.

### v3

@@ 3.2. General action estimates
Recalling (19), we first note that for every $\lambda\in(\lambda_{n},\lambda_{n+1})$ and $u\in Q_{\lambda}$, we have
@@ 4. Further estimates in the case of a metric measure space
Next we observe that, since $u\in Q_{\lambda}$, the function $u$ is an eigenfunction of the operator $A-V_{u}$ corresponding to the eigenvalue $\lambda$. Here, we note that $A-V_{u}$ is a self adjoint operator defined as a form sum of the operators $A$ and $V_{u}$. Here we note that, by Hölder’s inequality, $V_{u}$ is a well-defined and continuous multiplication operator $L^{p}(\Omega)\to L^{p^{\prime}}(\Omega)$. Therefore, by (38) and duality, it also defines a compact operator $E\to E^{\prime}$. In other words, $V_{u}$ is a symmetric and relatively form compact perturbation of the operator $A$, which implies that $A-V_{u}$ is also bounded from below and has a compact resolvent, see [22, Chapter XIII.4 and Problem XIII.39]. Therefore $A-V_{u}$ admits a sequence of eigenvalues
@@ 5.2. Weyl’s law, large spectral gaps and large $L^{\infty}$ norms
$V_{u}\in L^{\infty}(\mathcal{G};\mathbb{R}),\qquad V_{u}(x)=\left\{\begin{aligned} &\frac{f(u(x))}{u(x)},&&\qquad\text{$u(x)\neq 0$,}\\ &0,&&\qquad\text{$u(x)=0$}\end{aligned}\right$.
@@
(50) $\lambda_{n}(u)=\inf_{\stackrel{{\scriptstyle E^{\prime}\subset E}}{{\dim E^{\prime}\geq n}}}\sup_{v\in E^{\prime}\setminus\{0\}}\frac{1}{\|v\|_{H}^{2}}\Bigl(\langle A^{1/2}v,A^{1/2}v\rangle-\int_{\mathcal{G}}V_{u}v^{2}\,\mathrm{d}x\Bigr).$
@@
$\frac{1}{\|v\|_{H}^{2}}\Bigl|\int_{\mathcal{G}}V_{u}v^{2}\,\mathrm{d}x\Bigr|\leq\|V_{u}\|_{L^{\infty}(\mathcal{G})}\qquad\text{for every $u\in E$, $v\in E\setminus\{0\}$,}$
@@
$\lambda_{n+1}-\lambda\leq\lambda_{n+1}-\lambda_{n+1}(u)\leq\|V_{u}\|_{L^{\infty}(\mathcal{G})}\leq\tilde{g}(\|u\|_{L^{\infty}(\mathcal{G})}).$
Here, the last inequality follows from the pointwise bound $|V_{u}(x)|\leq\tilde{g}(|u(x)|)$ combined with the monotonicity of $\tilde{g}$. This proves the claim. ∎
@@ 5.3. ODE arguments and the $L^{\infty}$ to $L^{2}$ estimate
for all $M\geq R$. Since $\frac{1}{\kappa_{0}}R\geq s_{1}$, Lemma 5.7 implies that $\tau_{\lambda,M}\leq\ell/2$ for all $M\geq R$. Hence, the solution $u_{\mid e}$ completes at least two full periods on $e$. Let $I\subseteq e$ be a union of $\lfloor|e|/\tau_{\lambda,M}\rfloor$ consecutive full periods of $u_{\mid e}$, so that $|I|=\lfloor|e|/\tau_{\lambda,M}\rfloor\,\tau_{\lambda,M}\geq|e|-\tau_{\lambda,M}\geq|e|/2\geq\ell/2$. Applying Lemma 5.5 to each period, we obtain
@@ 5.4. Sign of solutions in the presence of a constant-sign eigenfunction
Taking scalar products with $\xi_{i}$, we obtain
$(Au,\xi_{i})_{L^{2}}=\lambda\int_{\mathcal{G}}u\xi_{i}\,\mathrm{d}x+\int_{\mathcal{G}}f(u)\xi_{i}\,\mathrm{d}x.$
Since $\xi_{i}$ is an eigenfunction of $A$, we have that
$(Au,\xi_{i})_{L^{2}}=\lambda_{i}\int_{\mathcal{G}}u\xi_{i}\,\mathrm{d}x.$
@@
$(\lambda_{i}-\lambda)\int_{\mathcal{G}}u\xi_{i}\,\mathrm{d}x=\int_{\mathcal{G}}f(u)\xi_{i}\,\mathrm{d}x,$
@@
As already pointed out in the introduction, the first eigenfunction is positive in $\mathcal{G}$ in the case of continuity-Kirchhoff conditions. For a further study of conditions ensuring positivity of the first eigenfunction, one may refer to [17, Section 4.5].
@@ 6. The nonlinear biharmonic equation on the $2$-torus
denotes the sums of two squares function. A result from analytic number theory (see [14, Chapter 16.9]) implies that
@@
where $d(j)$ is the number of divisors of $j$. From this inequality, one can deduce (see [14, Theorem 315]) that for every $\varepsilon>0$ there exists $C_{\varepsilon}>0$ with
@@ Acknowledgements
The authors thank Louis Jeanjean for valuable comments. D.G. is a Francqui Fellow of the Belgian American Educational Foundation at Brown University and was a Research Fellow of the F.R.S.-FNRS when this research was initiated. He would like to thank the Francqui Foundation, the Belgian American Educational Foundation, the University of Mons, the F.R.S.-FNRS and the Goethe-Universität Frankfurt for supporting this research.
Statements and declarations
The authors have no relevant financial or non-financial interests to disclose.
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.